# Vision-driven Autonomous Flight of UAV Along River Using Deep Reinforcement Learning with Dynamic Expert Guidance

### v1

@@ Abstract
Vision-driven autonomous flight and obstacle avoidance of Unmanned Aerial Vehicles (UAVs) along complex riverine environments for tasks like rescue and surveillance requires a robust control policy, which is yet difficult to obtain due to the shortage of trainable river environment simulators and reward sparsity in such environments. To easily verify the navigation controller performance for the river following task before real-world deployment, we developed a trainable photo-realistic dynamics-free riverine simulation environment using Unity. Successful river following trajectories in the environment are manually collected and Behavior Clone (BC) is used to train an Imitation Learning (IL) agent to mimic expert behavior and generate expert guidance. Finally, a framework is proposed to train a Deep Reinforcement Learning (DRL) agent using BC expert guidance and improve the expert policy online by sampling good demonstrations produced by the DRL to increase convergence rate and policy performance. This framework is able to solve the along-river autonomous navigation task and outperform baseline RL and IL methods. The code and trainable environments are available [^1] https://github.com/lijianwen1997/Deep-Reinforcement-Learning-with-Dynamic-Expert-Guidance.
@@ I INTRODUCTION
Autonomous flight of UAVs over inland waterways has been investigated in applications like water sampling [1], inspection of irrigation system [2], and cooperative drone-boat navigation tasks [3, 4, 5]. However, very few works addressed the along-river autonomous navigation tasks of UAVs, which can be applied in rescue, surveillance, and inspection missions over inland waterways. Due to seasonal changes of river channels’ geometry and water level, manually planned GPS waypoints may not suit for accurate river following missions. Besides, when navigating in cluttered riverine environments, UAVs also need to avoid obstacles like trees and bridges, while being above the water.
To address the above challenges, a camera is used as the major perception source for the UAV to take advantage of information embedded in images and the recent development in image processing. To reduce the cost and safety issues for real-world experiments, we developed a series of riverine simulation environments using Unity [6] to verify the feasibility of controlling a UAV to fly along a river with only photo-realistic vision input. The impact of environmental forces on UAV is not considered in the simulation environment and the UAV action in the 3D space is discretized to alleviate the navigation challenge of operating without a map along rivers.
Using the above simplification, we decouple the optimization of high-level local waypoint-based control policies or navigation planners for UAVs from mid-level velocity control, which is more suitable for the river following task due to its larger acceptable operation space compared to autonomous driving tasks for cars.
We pre-train an Imitation Learning (IL) agent using human demonstrations to provide basic guidance, then combine Reinforcement Learning (RL) algorithm with an online IL algorithm that accumulates and learns from decent agent experiences, to boost the learning speed and final agent performance. This will overcome the challenge of obtaining a performant controller using only RL algorithms since the only positive reward is given when the UAV traverses any newly visited discretized river segment. This training framework is examined to outperform Proximal Policy Optimization (PPO), Behavior Clone (BC), and PPO with static BC guidance.
@@
  • The creation of the trainable photo-realistic riverine simulation environment, which bridges the sim2real gap for vision-driven autonomous navigation of UAVs in the river following task.
  • The development of Deep Reinforcement Learning with a dynamic expert guidance framework that interactively trains Reinforcement Learning and Imitation Learning agents to reduce the training time and increase the overall performance.
@@ II RELATED WORK
Autonomous river following in a relatively simple curved river segment has been tested using a water segmentation mask from HSV threshed images [7, 8]. This image processing technique is preferable in terms of its implementation simplicity but has vulnerabilities like unverified generalization ability to rivers under different geometrical and light conditions, as well as the high safety hazards and experimental costs.
Game engines as simulation environments have brought more advantages for vision-driven Reinforcement Learning (RL) training, due to their flexibility in customizing environmental parameters, their low cost as a testbed before real-world deployment, and most importantly, their photo-realism to reduce the sim2real gap in air [9, 10], land [11, 12] (indoor/outdoor), and even underwater applications [13].
@@
To achieve a safer and more robust autonomous navigation controller in a diverse and adaptive riverine domain, researchers utilized Unreal Engine to train the agent for the river following task [14, 15]. Their proof of the superiority of the Variational Auto Encoder (VAE) [16] for image encoding, over Linear Regression method or end-to-end controller training, is adopted in our work. However, the reset criteria and the demonstrative trajectory collection approaches of their river simulation environment are manual-based, the agent action is limited to 2D planes, and the environment is not publicly available. To overcome these difficulties, our Unity-based river simulator can be automatically trained and tested for the river following task, supports 3D action space and collision detection, and is made public as an OpenAI gym [17] environment, thanks to Unity ML-Agents Toolkit [6].
Reinforcement learning(RL) and Imitation Learning(IL) are two powerful techniques with significant applications in the field of robotics. Robots can learn optimal behavior through trial and error through RL. Robots can learn to avoid obstacles [18], plan paths [19], and perform tasks like autonomous driving [20] through RL even in complicated, dynamic situations. On the other hand, IL [21] leverages expert demonstrations to train robots. By monitoring and imitating the actions of experts, robots can acquire complex skills rapidly. This is valuable in scenarios where human expertise is readily available, as it allows robots to learn efficiently without extensive programming [22].
However, both of these approaches have some limitations. RL frequently requires a large number of interactions with the environment to learn optimal policies, which can be time-consuming and expensive. Recently, Researchers have combined RL and IL to accelerate the training of RL using IL’s guidance. For example, [23] proposed to use the imitative expert policy to guide the learning of the DRL agent by regularizing the divergence between the DRL agent’s policy and the expert policy. [24] proposed a framework that uses a loss function that combines behavior cloning and 1-step Q-learning losses to enable the transition from BC to RL without performance degradation. However, after the DRL agent has developed a policy that is similar to the expert policy if the DRL agent tries to update the policy to improve the performance, it will be constrained by the expert policy or the value penalty thus the improvement is hindered.
On the other hand, as they have predominantly learned to imitate the demonstrations, robots trained through IL may not be able to adapt well to unexpected circumstances. DAgger [25] improves on BC by training on a dataset that better resembles the observations the expert policy is likely to encounter. However, the IL agent is also constrained by the expert. Thus to overcome these limitations, we propose an interactive framework called DRL with dynamic expert guidance that trains an RL agent using expert guidance and improves the expert policy online by sampling demonstrations produced by the DRL expert so that IL and RL can be improved jointly.
@@
(a) Comparison of image captured in Wildcat Creek, Indiana, USA (left) and image from Unity river environment (right) that are alike in overall components layout and texture appearance.
(b) Orthographic images of the training (left) and testing (right) river simulation environments.
Figure 1: Real and simulation river environments.
@@ III SIMULATION
III-A Photo-realistic River Environments
The riverine simulation environment is built in Unity using River Auto Material 2019  [^2] https://assetstore.unity.com/packages/tools/terrain/r-a-m-2019-river-auto-material-2019-145937 and Mountain Environment  [^3] https://assetstore.unity.com/packages/3d/vegetation/mountain-environment-dynamic-nature-191851. The river spline tool is provided by the former package to support fast, flexible and scriptable river building with various realistic water textures and shaders. The latter package provides a huge library of scanned forest assets like trees, branches, logs, and rocks, which helps to enrich the biomes of the riverine scene, as well as makes it more alike the real-world river scene components and appearance, as shown in Figure 0(a).
Aside from the realistic river and terrain trees, the riverine simulation environment features several additional aspects that not only mimic the general objects in real-world rivers, creeks, and streams, but also make the training of river following agents more difficult: 1) river channels with varied widths and depths, 2) river tributaries that diverge from the main water stream, 3) bridges over the river that can obstruct the navigable path of UAVs, and 4) sharp and slow turns. (Figure 2)
Several environment design choices are made for better photo-realism and more robust training. Instead of continuous velocity control, discrete position control for UAV movement is used, which eliminates the efforts of controller dynamics adaptation in real-world riverine environments with various wind speeds, water surface effects, etc. Thus we will name the UAV agent as the camera agent in the contents hereafter. The circular shape of rivers in our simulated environments is designed for better camera view coherence without seeing too many unrealistic un-modeled parts (sky, terrains, etc.) as the camera agent moves to any valid pose along the river. Besides, in the circular river environment, no matter where the camera agent is reset at the beginning of a training episode, the agent policy always has a chance to learn to traverse the whole river before ending this episode, making the collected trajectory lengths more uniform and independent of agent reset positions.
(a) Varying widths and depths
(b) Tributary
(c) Bridge
(d) Challenging U turns
Figure 2: Key features of the Unity river simulation environment.
TABLE I: Multi-discrete action values of the camera agent in the actual execution order, whose controlling logic is the same as remote controllers of common UAVs.
@@ III-A Photo-realistic River Environments
III-B Camera Agent
@@
The simulation agent, created using Unity ML-Agents Toolkit [6], is a camera that looks forward with a fixed pitch angle (20 degrees in our experiments) with respect to the drone frame. Since no dynamics is included in this simulation, and all control actions are discrete, the drone camera itself is viewed as the agent. The camera agent is controlled by a four-dimensional multi-discrete action that maps from the command inputs of the common remote controller of UAVs: 1) translation up and down - throttle, 2) horizontal rotation counter-clockwise and clockwise - yaw rate, 3) translation forward and backward - pitch, 4) translation leftward and rightward - roll. Each action branch has three discrete values $\{0,1,2\}$, where $1$ means no action on that branch. The actions are applied in the drone coordinate frame (instead of the camera frame) with the right-hand rule, where the positive X axis points along the drone head and the positive Z axis points upward to the sky. The specific values of movement caused by actions are in Table I.
III-C Gym Environment
@@
A river can be abstracted as a 3D Catmull-Rom spline [26] that is interpolated from several control points. The river surface is formed if we stretch this spline bi-directionally along some direction. This closed spline is lifted up to form a 3D bounding volume to limit the height of the camera agent, as well as its horizontal position. The tangents along the spline provide failure judgment when the camera agent’s yaw angle is far from either tangent angle of the nearest point on this spline with the projected agent position. This prohibits the turnaround motion of the agent since the built river is circular so that the agent can finish the whole-river navigation without any turnaround. Even if the agent does not trigger the above two ”done” criteria, lingering around or doing continuous meaningless movements with no reward improvement (over 50 steps in our experiments) will also trigger environment reset, and so does collision with bridges. If the environment ”done” is triggered during training, the camera agent will be reset at a random point inside the spline-based bounding volume, with a random yaw near either spline tangent of the nearest spline point. The above criteria for the river following task can be better visualized in Figure 5.
@@
Figure 3: Demonstrative diagram of valid activity space and acceptable yaw range of camera agent in the river following task. $h_{1}=1m,h_{2}=15m,\alpha=150^{\circ}$ in our experiments. The drone picture is from  [^5] https://www.jaycar.com.au/r-c-fpv-drone-with-1080p-camera/p/GT4118.
@@
The river central spline can be simplified as a sequence of connected line segments with nearly the same length. To have a decent baseline RL algorithm to compare, and to intuitively sense the learning capabilities of different learning algorithms, shaped reward is given to the camera agent if the projection of its positions on the spline before and after an action covers any un-visited line segments. The line segment where the camera agent’s current position projects onto is denoted as visited. If the camera agent has visited all the river splines, it will get the highest reward, then the environment will be reset. For collision detection, a mesh collider is used for bridges, and a box collider with a side length of $0.5m$ is used for the camera agent. The overall reward function for the autonomous river following task is given below.
@@
$r(s,a)=\begin{cases}-1,&\text{agent out of spline bounding volume}\\ -1,&\text{agent yaw diverts too much}\\ -1,&\text{agent collides with bridge}\\ -1,&\text{agent no progress over certain steps}\\ r_{shaped}&\text{agent covers un-visited spline segment}\\ 0,&\text{otherwise}\end{cases}$ (1)
@@
where $r_{shaped}$ is the percentage of newly-visited segments $n$ after the action $a$ over all segments number $N$, multiplied by a constant value, $10$ in this case.
$r_{shaped}=10*n/N\$ (2)
All negative reward conditions denote the reset of the environment (Equation 1). Besides, no final positive reward is given when the camera agent has visited all river spline segments since the amplified shaped reward, Equation 2, already gives informative guidance for the river following task. Due to the same reason, no environment reset is applied for backup movement or staying still of the agent.

### v2

# Synergistic Reinforcement and Imitation Learning for Vision-driven Autonomous Flight of UAV Along River

@@ Abstract
Vision-driven autonomous flight and obstacle avoidance of Unmanned Aerial Vehicles (UAVs) along complex riverine environments for tasks like rescue and surveillance requires a robust control policy, which is yet difficult to obtain due to the shortage of trainable riverine environment simulators. To easily verify the vision-based navigation controller performance for the river following task before real-world deployment, we developed a trainable photo-realistic dynamics-free riverine simulation environment using Unity. In this paper, we address the shortcomings that vanilla Reinforcement Learning (RL) algorithm encounters in learning a navigation policy within this partially observable, non-Markovian environment. We propose a synergistic approach that integrates RL and Imitation Learning (IL). Initially, an IL expert is trained on manually collected demonstrations, which then guides the RL policy training process. Concurrently, experiences generated by the RL agent are utilized to re-train the IL expert, enhancing its ability to generalize to unseen data. By leveraging the strengths of both RL and IL, this framework achieves a faster convergence rate and higher performance compared to pure RL, pure IL, and RL combined with static IL algorithms. The results validate the efficacy of the proposed method in terms of both task completion and efficiency. The code and trainable environments are available [^1] https://github.com/lijianwen1997/Synergistic-Reinforcement-and-Imitation-Learning.
@@ I INTRODUCTION
Autonomous flight of UAVs over inland waterways has been investigated in applications like water sampling [1], inspection of irrigation system [2], and cooperative drone-boat navigation tasks [3, 4, 5]. However, few studies addressed the along-river autonomous navigation tasks of UAVs, which can be applied in rescue, surveillance, and inspection missions over inland waterways. Seasonal variations in river channel geometry and water levels render manually planned GPS waypoints less reliable for precise river following missions. Moreover, in cluttered riverine environments, UAVs must navigate while avoiding obstacles such as trees and bridges, all while maintaining a safe altitude above the water.
A camera can be used as the major perception source for the UAV to take advantage of information embedded in images and the recent developments in image processing. To reduce the cost and safety issues for real-world experiments, we developed a riverine simulation environment using Unity [6] to verify the feasibility of controlling a UAV to fly along a river with only photo-realistic vision input.
Figure 1: System architecture. A human expert collects good trajectories before training and the transitions are represented by blue arrows. $a^{H}_{t}$ denotes the human expert’s action. The transitions during training are represented by black arrows. $a^{E}_{t}$ denotes the IL expert’s action.
River following can be classified as a track following task, which poses significant training challenges as the environment does not naturally conform to a Markov Decision Process (MDP) due to the agent’s partial observability and the non-Markovian reward function [7]. Specifically, the first-person camera view of the UAV agent in the riverine scene may lack sufficient water pixels to guide the agent along the river, or it may contain bifurcated river channels, confusing the agent’s navigation decisions. Moreover, the UAV agent receives positive rewards only upon visiting new river segments, with no reward for idling above already visited segments. Consequently, the immediate reward depends on both past observations and actions, rather than solely on the current ones, rendering the task a Partially Observable Markov Decision Process (POMDP) with non-Markovian rewards. To address this challenge, combined approaches of RL and Learning from Demonstration (LfD) have been proposed [8, 9, 10]. However, how to effectively leverage the complementary advantages of imitative learning from demonstrations and reinforcement learning from experience for the track following task remains an area requiring further research and exploration.
To investigate into this problem, we pre-train an IL expert using human demonstrations to provide basic guidance, then combine a RL algorithm with the online-updating IL algorithm that accumulates and learns from decent agent experiences, to mimic the progressive expert behaviors while learning. In our experiments, Proximal Policy Optimization (PPO [11]) is chosen as the RL agent, Behavior Cloning (BC [12]) the expert. This training framework is examined to boost the learning speed and improve final agent performance in the track following tasks, outperforming PPO, StaticBC, DynamicBC and PPO with StaticBC guidance baselines.
@@
  • The development of synergistic reinforcement and imitation learning framework that interactively trains RL agent and IL expert to reduce the convergence time and increase the overall performance.
  • The creation of the trainable photo-realistic riverine simulation environment, which is publicly available, and serves as a testbed for vision-driven autonomous navigation of UAVs in the river following task.
@@ II RELATED WORK
Autonomous river following in a relatively simple curved river segment has been tested using a water segmentation mask from HSV threshed images [13, 14]. This water mask extraction technique and the control method both require careful manual tuning, especially for complex riverine environments with obstacles and multiple river channels. To alleviate laborsome manual tuning of the river following controller, modern image embedding methods and RL algorithms can be investigated in a photo-realistic riverine simulator.
@@
Game engines as simulation environments have brought more advantages for vision-driven RL training, due to their flexibility in customizing environmental parameters, their low cost as a testbed before real-world deployment, and most importantly, their photo-realism to reduce the Sim2Real gap in air [15, 16], land [17, 18] (indoor/outdoor), and even underwater applications [19]. To achieve a safer and more robust autonomous navigation controller in a diverse and adaptive riverine domain, researchers utilized Unreal Engine to train the agent for the river following task [20, 21]. Their proof of the superiority of the Variational Auto Encoder (VAE) [22] for image encoding, over end-to-end controller training, is adopted in our work. However, the reset criteria of their river simulation environment are manual-based, the agent action is limited to a 2D plane, and the environment is not publicly available. To overcome these difficulties, our Unity-based riverine simulator can be automatically trained and tested for the river following task, supports 3D action space and collision detection, and is made public as an OpenAI gym [23] environment, thanks to Unity ML-Agents Toolkit [6].
RL and IL are potent techniques in robotics. RL enables robots to learn optimal behavior through trial and error, mastering tasks like obstacle avoidance [24], path planning [25], and autonomous driving [26] even in dynamic environments. Conversely, IL harnesses expert demonstrations to swiftly train robots, allowing them to mimic complex skills. This method is particularly valuable when human expertise is accessible, streamlining robot learning without extensive programming [27]. However, both approaches have limitations. RL training can be time-consuming and costly, often demanding a substantial number of interactions with the environment to learn optimal policies. IL’s effectiveness is constrained by its reliance on static training datasets, making it challenging to generalize to novel scenarios.
To address these issues, researchers have begun integrating RL and IL to enhance training efficiency and performance. For instance, DAgger [28] enhances BC by training on a dynamically updating dataset mirroring the observations the trained policy is expected to encounter. However, it necessitates continual online querying of optimal actions from a human expert. Other approaches include frameworks that use the guidance of IL to train RL [29, 30], regulating the discrepancy between the agent’s policy and the expert’s. In particular, [30] combines BC and 1-step Q-learning losses to smoothly transition from IL to RL without performance degradation. However, for these methods, the IL policy is static, and further improvements of RL agents are impeded by the constraints imposed by the expert policy.
To overcome these limitations, we introduce an interactive framework called ”Synergistic Reinforcement and Imitation Learning.” This framework trains an RL agent using IL expert advice while simultaneously refining the expert policy through online sampling of high-quality demonstrations generated by the RL agent. By facilitating collaborative improvement between IL and RL methodologies, this approach enhances training efficiency and overall performance.
@@
(a) Comparison of an image captured in Wildcat Creek, Indiana, USA (left) and image from Unity river environment (right) that are alike in overall components layout and texture appearance.
(b) Orthographic images of the riverine simulation environment and the CliffCircular environment.
Figure 2: Real and simulated riverine environments, and CliffCircular environment (with two extra random cliff grids) for river/track following task.
@@ III SIMULATION
III-A Photo-realistic Riverine Environment
The riverine simulation environment is built in Unity using River Auto Material 2019 and Mountain Environment . The river spline tool is provided by the former package to support fast, flexible, and scriptable river building with various realistic water textures and shaders. The latter package provides a huge library of scanned forest assets like trees, branches, logs, and rocks, which helps to enrich the biomes of the riverine scene, as well as makes it more alike the real-world riverine scene components and appearance, as shown in Fig. 2(a).
Aside from the realistic river and terrain trees, the riverine simulation environment features several additional aspects that not only mimic the general objects in real-world rivers, creeks, and streams, but also make the training of river following agents more difficult: 1) river channels with varied widths and depths, 2) river tributaries that diverge from the main water stream, 3) bridges over the river that can obstruct the navigable path of UAVs, and 4) sharp and slow turns.
Several environment design choices are made for better photo-realism and more robust training. Instead of continuous velocity control, discrete waypoint control for UAV movement is used, which eliminates the efforts of controller dynamics adaptation in real-world riverine environments with various wind speeds, water surface effects, etc. Thus we will name the UAV agent as the camera agent in the contents hereafter. The circular shape of the river in our simulated environment is designed for better camera view coherence without seeing too many unrealistic un-modeled parts (sky, terrains, etc.) as the camera agent moves to any valid pose along the river. The annulus design of the riverine environment guarantees consistent episodic rewards achievable regardless of the initial reset position of the camera agent at the beginning of each episode.
TABLE I: Multi-discrete action values of the camera agent.
@@ III-A Photo-realistic Riverine Environment
III-B Riverine Camera Agent
The simulation agent, created using Unity ML-Agents Toolkit [6], is a camera that looks forward with a fixed pitch angle ($20$ degrees in our experiments) with respect to the drone frame.
@@ III-B Riverine Camera Agent
The camera agent is controlled by a four-dimensional multi-discrete action that maps from the command inputs of the common remote controller of UAVs: 1) translation up and down - throttle, 2) horizontal rotation counter-clockwise and clockwise - yaw rate, 3) translation forward and backward - pitch, 4) translation leftward and rightward - roll. Each action branch has three discrete values $\{0,1,2\}$, where $1$ means no action on that branch. The specific values of movement caused by actions are in Table I.
III-C Riverine Gym Environment
@@
A river can be abstracted as a 3D Catmull-Rom spline [31] that is interpolated from several control points. The river surface is formed if we stretch this spline bi-directionally along some direction. This closed spline is lifted up to form a 3D bounding volume to limit the height of the camera agent, as well as its horizontal position. The tangents along the spline provide failure judgment when the camera agent’s yaw angle is far from either tangent angle of the nearest point on this spline with the projected agent position. This prohibits the turnaround motion of the agent since the built river is circular so that the agent can finish the whole-river navigation without any turnaround. Lingering around or doing continuous meaningless movements with no reward improvement (over $50$ steps in our experiments) will also trigger environment reset, and so does collision with bridges. If the environment ”done” is triggered during training, the camera agent will be reset at a random point inside the spline-based bounding volume, with a random yaw near either spline tangent of the nearest spline point. The above criteria for the river following task can be better visualized in Fig. 3.
@@
Figure 3: Demonstrative diagram of valid activity space and acceptable yaw range of camera agent in the river following task. $h_{1}=1m,h_{2}=15m,\alpha=150^{\circ}$ in our experiments.
@@
The river central spline can be simplified as a sequence of connected line segments with nearly the same length. To have a decent baseline RL algorithm to compare, and to intuitively sense the learning capabilities of different learning algorithms, shaped reward is given to the camera agent if the projection of its positions on the spline before and after an action covers any un-visited line segments. The line segment where the camera agent’s current position projects onto is denoted as visited. If the camera agent has visited all the river spline segments, it will get the highest reward, then the environment will be reset. For collision detection, a mesh collider is used for bridges, and a box collider with a side length of $0.5$m is used for the camera agent. The overall reward function for the autonomous river following task is given below.
@@
$r(s,a)=\begin{cases}-1,&\text{agent fails}\\ r_{shaped}&\text{agent covers un-visited spline segment}\\ 0,&\text{otherwise}\end{cases}$ (1)
@@
Agent fails when 1) out of spline bounding volume, 2) yaw diverts too much, 3) collides with bridge, and 4) no progress over certain steps. $r_{shaped}$ is the percentage of newly-visited segments $n$ after the action $a$ over all segments number $N$, multiplied by a constant value, $10$ in this case: $r_{shaped}=10*n/N$. All negative reward conditions denote the reset of the environment (Eq. 1).
III-D CliffCircular Environment
CliffCircular (Fig. 2(b)) is derived from the CliffWalking environment [32, 33] and tailored for the track following task. Like the camera agent in the riverine environment, the agent in CliffCircular must navigate along walkable grids without stepping onto cliffs until it circumnavigates the central cliff block. Additionally, CliffCircular presents partial observability and non-Markovian characteristics due to the local observation of cliffs surrounding the agent’s current grid, the flexibility in determining the circulation direction, and the nature of the track following task, which necessitates consideration of historical information.
Figure 4: CliffCircular environment diagram. Central track grids, agent location and observation grids are laid on top of the color layers of obstacle and walkable grids.
CliffCircular is a $12$x12121212 grid world, illustrated in Fig. 4. The observation space is the flattened $5$x5555 grids around the agent indicating cliff. Action space includes five discrete operations: no operation, up, right, down and left. The $+1$ reward will be given if agent’s projection on the central track grid is newly visited, $-100$ if agent falls off cliff, otherwise 00. Episode will be reset if cliff is triggered or all central track grids are visited. At the beginning of each episode, the agent’s location and two additional cliff grids are randomly initialized. This setup encourages the policy to learn a safe and robust track following strategy. A TimeLimit environment wrapper of length $128$ is used.